OpenStudy (anonymous):

G(x)= 4Sin((pi)(x))-3 Solve: ((Pi)(x))-3=0 and ((Pi)(x))-3 = (2)(PI) Please help.

OpenStudy (campbell_st):

well start with the basics add 3 to both sides of the equation next divide both sides of the equation by 4 find sin^-1 of (3/4) then divide by pi I'll assume you're working in radians

OpenStudy (anonymous):

I'm trying to find the interval of the sin graph using the two bottom equations

OpenStudy (anonymous):

I got 3/pi and 5 as my two answers. I don't know if they're correct.

OpenStudy (campbell_st):

well the way I read the question, you are looking for an angle so \[\pi \times x = \sin^{-1}(\frac{3}{4})\] so \[\pi \times x = 0.8481r\] and then solve for x from there

OpenStudy (anonymous):

The question asks me to graph this equation on the same graph as F(x)=4sin (pi)(x)

OpenStudy (anonymous):

I know that the amplitude is 4 and that the period is 2 but i don"t know the intervals to graph the G(x) equation

OpenStudy (campbell_st):

the amplitude is 4, and the period 2, the curve is centred on y = -3 so the range is 1 to -5 so the solution above gives the 1st quadrant value in radians where the curve is equal to zero. There is also a solution in the 2nd quadrant. the angle can be found by using \[(\pi - x) ... radians\] hope this helps.